John invested $20,000 in two different accounts, one paying 7% and the other 9%. If total interest earned for the year was $1,550, how much was invested in each account.
Answer by Earlsdon(6294)

Let x = the amount invested at 9% then (20,000-x) would be the amount invested at 7%. The interest on these two amounts can be expressed as:
0.09x+0.07(20000-x) = $1,550 Simplify this and solve for x.
0.09x+1400-0.07x = 1550
0.02x+1400 = 1550
0.02x = 150 then...
x = 7,500 and 20,000-x = $12,500
$7,500 was invested at 9% and $12,500 was invested at 7%